# Mutual benefits of social learning and algorithmic mediation for cumulative culture

### v1

@@ 3 Results
If not indicated otherwise, we present results for populations of $N=100$ agents that are embedded in a random social network, where each pair of individuals is connected with probability $p$ and thus has on average $k=pN$ neighbors. The effort budget of agents is set to $B=1000$, the costs for acquiring cultural traits are fixed at $C_{i}=10$, $C_{s}=5$, and $C_{r}=1$, and we assume $X=100$ cultural branches.
Due to the finite effort budget $B$ and the low probability of the death-birth process $q$, the system reaches a quasi-stationary state regardless of the other parameters. In the following, we will thus focus on these ”maximally achievable values”, i.e., population mean values at the stationary state. For brevity, we will refer to these values as ”mean group values” and mark them with a subscript max. Each simulation continues for $t=20N$ time steps and the last $200$ time steps are used to compute the mean group values. Our results are then reported as averages over $200$ realizations.
@@ 3.1 Network-based social learning
Figure 2 depicts the results for fully connected networks ($p=1$). Similar to the findings reported in [26], we observe that after the initial growth of $\bar{Z}$ the dynamics reach a quasi-stationary state where the mean group payoff approaches its maximum value, denoted as $\bar{Z}_{\mathrm{max}}$. In line with classical results in cultural evolution [12], we find that the maximum cultural complexity increases with the system size uncovering a logarithmic relationship. This relationship becomes more clear by plotting $\bar{Z}_{\mathrm{max}}$ as a function of $\log(N)$ (system size), as depicted in the right panel of Figure 2.
@@
Figure 3 shows the results for random networks with different connectivity, as parameterized by the average number of neighbours $k$ of an agent. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of cultural traits $\bar{T}_{\mathrm{max}}$, mean group payoff $\bar{Z}_{\mathrm{max}}$. The mean group level grows with the number of neighbors until it reaches a plateau for very dense networks. Instead, the number of discovered items $\bar{T}_{\mathrm{max}}$ decreases rapidly and monotonically with $k$. The most striking behavior is observed for the mean group payoff. Initially, for very sparse networks (low values of $k$) we observe an increase of $\bar{Z}_{\mathrm{max}}$with $k$, but after a maximum value is reached, for intermediate levels of connectivity, we observe that $\bar{Z}_{\mathrm{max}}$ decreases with $k$.
Figure 3: System behavior in a stationary state as a function of network connectivity $k$. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $X=100$
@@ 3.2 Algorithmic mediation
Next, we focus on the quantitative and qualitative impact of the algorithmic mediation on cumulative cultural evolution. The results are shown in Fig. 4. Only the mean group level, $\bar{L}_{\mathrm{max}}$, shows a monotonous increase with increasing $r$ for every value of $k$, see left panel of Fig. 4. In line with the observations from the previous section, a higher level of development is observed for denser networks. By contrast, the mean group payoff $\bar{Z}_{\mathrm{max}}$ gives rise to maximum values for intermediate levels of $r$, and these maximum values decrease with $k$. This suggests a trade-off between network connectivity and the ratio of recommendations, which can be seen more clearly in Fig. 5. The highest payoff is observed for very sparse networks (every agent has only one neighbor on average) and relatively high levels of recommendations ($r\approx 0.7$), which might be due to the high number of innovated traits $\bar{T}_{\mathrm{max}}$, for low values of $k$, as can be seen from the middle panel of Fig.  4). With fewer neighbors an individual benefits from a higher number of recommendations, but this number decreases with $k$. Note that both limiting cases of $r=0$ and $r=1$ never give rise to maximum values, independently of the connectivity, suggesting that network-based social learning and algorithmic-mediation of information are mutually beneficial.
@@
Figure 4: Stationary states for mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$ (from left to right) as a function of recommendation probability $r$, and for different network connectivities $k$ including complete graphs (CG). The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
@@
Figure 5: Heatmap of the mean group payoff $\bar{Z}_{\mathrm{max}}$ as a function of the average connectivity $k$ and probability of reccomendation $r$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
@@
From our numerical simulations, it appears that both the size of the system ($N$) and the number of possible development paths $X$, do not qualitatively affect our modeling results. We also did not observe non-trivial relationships between the parameters related to learning costs (i.e. $C_{i}$, $C_{s}$, $C_{r}$, $B$) and the measures of cultural accumulation. Intuitively, the more we increase the cost of learning or innovation, the lower the payoff of the group. Similarly, a larger effort budget translates into higher cultural accumulation, where the resulting dynamics depend on the network topology and the recommendation probability $r$. Nevertheless, our findings are applicable within the context of reasonable parameter values, where the costs of acquiring items are significantly lower than the effort budget ($C\ll B$). Finally, it is important to point out that although we focused on cases in which the costs of social learning are higher than the costs of algorithmic mediation, we confirmed that our results are robust for $C_{s}=C_{r}$.
@@ 4 Discussion
We investigated the effects of network-based social learning and algorithmic mediation on cultural accumulation by building on, and extending, previous modeling efforts from the field of cultural evolution [26]. In the first part, we considered the model without algorithmic mediation and found a strong dependence of cultural accumulation on population size. In particular, for fully connected networks, the ability of a population to accumulate culture showed a positive relation (logarithmic growth) with population size, which is consistent with established findings in cultural evolution [6, 35, 36, 37, 38, 39]. We then investigated cultural accumulation in random networks that differ in the density of connections, as quantified by the average number of neighbors $k$ of a single agent. Although increasing the link density always led to higher levels of development ($\bar{L}_{\mathrm{max}}$), this was not the case for the mean group payoff ($\bar{Z}_{\mathrm{max}}$) and the number of innovated traits ($\bar{T}_{\mathrm{max}}$), which peak at intermediate values of $k$ and monotonously decrease with $k$, respectively. These non-trivial behaviors indicate that cultural accumulation is subject to an exploration-exploitation trade-off between innovation (exploration) and social learning (exploitation), a phenomenon that was previously also observed in both experimental and theoretical works on social dynamics including cultural evolution [13, 28], wisdom of the crowds [40], as well as collective problem-solving [41, 31, 42, 43]. Dense random networks (i.e. networks with high values of $k$) generally allow to disseminate information about cultural traits efficiently (exploitation) and therefore lead to high cultural complexity, however, do so at the expense of cultural diversity (number of traits) as the rate of innovation is impaired due to an over-reliance on social learning. Furthermore, previous empirical and theoretical findings in cultural evolution and social learning studies suggest that collective performance depends on both the complexity of the task and the underlying network structure. In particular, it was found that performance in dense networks is often decreased for complex tasks [31, 42, 43], a phenomenon which is best explained by sufficient levels of transient diversity in less densely connected systems [44]. In Fig. 3, the shapes of $\bar{L}_{\mathrm{max}}$ and $\bar{Z}_{\mathrm{max}}$, as a function of $k$, therefore indicate that the different sub-processes of cultural accumulation vary in complexity: While reaching a high level of complexity is a rather simple problem (high value of $\bar{L}_{\mathrm{max}}$ in dense networks), it is more complex to collectively coordinate on high average payoffs (low values of $\bar{Z}_{\mathrm{max}}$ in dense networks). In other words, populations easily reach high level of development, but do not necessarily do so with high collective payoffs. In our model, collective payoffs are optimal at a medium level of connectivity, where agents optimally solve the explore-exploit trade-off, i.e., have access to the right amount of information through social learning while maintaining sufficient incentives to innovate.
@@
In the second part, we investigated the combined impact of network-based social learning and algorithmic mediation on cultural accumulation. For a fixed value of $k$, we find that increasing $r$—the ratio between algorithmic mediation over network-based social learning—strongly affects the overall dynamics: while the developmental level monotonously increases with $r$, the mean payoff gives rise to maximum values for intermediate levels of algorithmic mediation, which suggests a non-trivial interaction between $k$ and $r$. The less we are connected, the more we can benefit from algorithmic mediation. For a given link density, however, the benefits of increased information sharing via algorithmic mediation become outweighed by their detrimental consequences: too high levels of $r$ create too much of effective connectivity which decreases performance as $r\rightarrow 1$. This is in line, with the strongly decreasing number of traits as $r$ increases, indicating low levels of innovation. These results make a clear case for studying the dynamics of algorithmic mediation not in isolation but in relation to social learning strategies motivated by cognitive science (e.g. payoff-biased learning). In our model, both algorithmic mediation and social learning acting alone consistently lead to worse payoffs and lesser cultural traits than a combination of both processes. This behavior may point toward a general trade-off in socio-technical systems between access to information (and thus the possibility to exchange it) and the ability of individuals to innovate: while algorithms can be beneficial for cultural accumulation, they may also lead to detrimental effects if used excessively [45, 46].
Our research takes a first step in modeling the impact of personalized algorithmic mediation on cumulative cultural evolution, however, it does not come without limitations. By focusing on the algorithmic mediation of information, inspired by recommender systems or feed algorithms, we have not considered other types of intelligent algorithms that are increasingly intertwined with human culture, in particular, generative artificial intelligence [47]. Importantly, generative AI not only re-distributes information, but is able to create novel cultural variation, which is not captured in our models. Furthermore, in line with previous models, we considered a very simplistic cultural space, namely cultural branches that are independent from each other. A more realistic approach could therefore take into account that cultural innovations often arise through the recombination of artefacts from distant domains, i.e. modeling interdependencies between branches. Similarly, by investigating random networks and agents with identical social learning strategies we have omitted many of the behavioral and structural heterogeneities observed in real social systems [48, 49]. The use of more complex network models—such as scale-free [50] and small-world networks [51]—as well as agents with individual preferences are promising avenues for future research directions. Finally, we have operationalised algorithmic mediation as a simple system that provides agents with ”ideal” personalised recommendations. Therefore, extending our approach to more realistic models of algorithmic mediation, e.g. recommendations based on collaborative filtering [52], is a promising next step. Overall, our work is only a first step towards a better theoretical understanding of the impact of algorithms on cumulative cultural evolution. Ultimately, the predictions of our model must stand up to empirical testing, both through controlled laboratory experiments and real-world observational studies across different online environments.
Code availability
Simulation code is available at https://github.com/aga-cz/Mutual-benefits-of-SL-and-RS

### v2

@@ 3 Results
If not indicated otherwise, we present results for populations of $N=100$ agents that are embedded in a random social network, where each pair of individuals is connected with probability $p$. Hence, each agent has on average $k=p(N-1)$ neighbours. The effort budget of agents is set to $B=1000$, the costs for acquiring cultural traits are fixed at $C_{i}=10$, $C_{s}=5$, and $C_{r}=1$, and we assume $X=100$ cultural branches. The SM contains a sensitivity analysis of the model with respect to these parameters.
Due to the finite effort budget $B$ and the low probability of the death-birth process $q$, the system reaches a quasi-stationary state regardless of the remaining parameters. In the following, we will thus focus on ”maximally achievable values”, i.e., population mean values at the stationary state for performance ($\bar{Z}$), diversity ($\bar{T}$), and complexity ($\bar{L}$). For brevity, we will refer to these values as ”mean group values” and mark them with a subscript max. Each simulation continues for $t=20N$ time steps and the last $200$ time steps are used to compute the mean group values. Our results are reported as averages over $200$ realizations.
@@ 3.1 Network-based social learning
Figure 2 depicts the results for fully connected networks ($p=1$). Similar to the findings reported in [31], we observe that after the initial growth of $\bar{Z}$ the dynamics reach a quasi-stationary state where the mean group payoff approaches its maximum value, denoted as $\bar{Z}_{\mathrm{max}}$. In line with classical results in cultural evolution [13], we find that the maximum group performance increases with the system size uncovering a logarithmic relationship. This relationship becomes more clear by plotting $\bar{Z}_{\mathrm{max}}$ as a function of $\log(N)$ (system size), as depicted in the right panel of Figure 2.
@@
Figure 3 shows the results for random networks with different connectivity, as parameterized by the average number of neighbours $k$ of an agent. From left to right we depict the mean group level $\bar{L}_{\mathrm{max}}$, the mean number of cultural traits $\bar{T}_{\mathrm{max}}$, and the mean group payoff $\bar{Z}_{\mathrm{max}}$ as functions of $k$. The mean group level grows with the number of neighbours until it reaches a plateau for very dense networks. Instead, the number of discovered items $\bar{T}_{\mathrm{max}}$ decreases rapidly and monotonically with $k$. The most striking behavior is observed for the mean group payoff. Initially, for very sparse networks (low values of $k$) we observe an increase of $\bar{Z}_{\mathrm{max}}$with $k$, but after a maximum value is reached, for intermediate levels of connectivity, we observe that $\bar{Z}_{\mathrm{max}}$ decreases with $k$.
Figure 3: Stationary states as functions of network connectivity $k$. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $X=100$
@@ 3.2 Algorithmic mediation
In Fig. 4, we show the results for the impact of algorithmic mediation on cumulative cultural evolution. Only the mean group level, $\bar{L}_{\mathrm{max}}$, shows a monotonous increase with increasing ratio of algorithmic mediation $r$ for every value of $k$, see left panel of Fig. 4. In line with the observations from the previous section, a higher level of development is observed for denser networks. By contrast, the mean group payoff $\bar{Z}_{\mathrm{max}}$ gives rise to maximum values for intermediate levels of $r$, and these maximum values decrease with $k$. This suggests a trade-off between the network connectivity ($k$) and the ratio of algorithmic mediation ($r$), which can be seen more clearly in Fig. 5. The highest payoff is observed for very sparse networks (every agent has only one neighbour on average) and relatively high levels of algorithmic mediation ($r\approx 0.7$), which might be due to the high number of innovated traits $\bar{T}_{\mathrm{max}}$, for low values of $k$, as can be seen from the bottom panel of Fig.  4. With fewer neighbours an individual benefits from a higher level of algorithmic mediation, but this trend decreases with $k$. Note that both limiting cases of $r=0$ and $r=1$ never give rise to maximum values, independently of the connectivity, suggesting that network-based social learning and algorithmic-mediation of information are mutually beneficial.
@@
Figure 4: Stationary states for mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$ (from left to right) as a function of the ratio of algorithmic mediation $r$, and for different network connectivities $k$ including complete graphs (CG). The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
@@
Figure 5: Heatmap of the mean group payoff $\bar{Z}_{\mathrm{max}}$ as a function of the average connectivity $k$ and the ratio of algorithmic mediation $r$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$.
@@
From our numerical simulations, it appears that both the size of the system ($N$) and the number of possible development paths ($X$), do not qualitatively affect our modeling results. We also did not observe non-trivial relationships between the effort budget ($B$) and the measures of cultural accumulation. Intuitively, a larger effort budget translates into higher payoffs, where the resulting dynamics depend on the network topology and the ratio of algorithmic mediation $r$. In the SM, we report additional simulation results on how the costs of acquiring cultural traits (i.e. $C_{i}$, $C_{s}$, $C_{r}$) affect the dynamics. We find that lower values of the learning costs are associated with higher measures of cultural accumulation, where the impact of ($C_{i}$, $C_{s}$, $C_{r}$) is most pronounced for the mean group level, and much smaller for the mean number of traits and the mean group payoff. We also noticed some qualitative changes of $\bar{Z}_{max}$ as a function of the ratio of algorithmic mediation $r$ for different values of $C_{i}$, $C_{s}$, and $C_{r}$. Importantly, we find that in every investigated case our main result remains unchanged, i.e. intermediate values of $r$ give rise to the maximum performance (peak values of $\bar{Z}_{max}$), where the exact location of that optimum depends on the costs of innovation, social learning and algorithmic mediation, as reported in the SM.
@@ 4 Discussion
We investigated the effects of network-based social learning and algorithmic mediation on cultural accumulation by building on, and extending, previous modeling efforts in cultural evolution [31]. In the first part, we considered the model without algorithmic mediation and found a strong dependence of cultural accumulation on population size. In particular, for fully connected networks, the population’s ability to accumulate culture exhibited a positive relationship with population size, characterized by logarithmic growth. This result aligns with established findings in cultural evolution, highlighting the role of population size in fostering cumulative cultural dynamics [6, 40, 41, 42, 43, 44]. We then investigated cultural accumulation in random networks that differ in the density of connections, as quantified by the average number of neighbours $k$ of a single agent. Although increasing the link density always led to higher levels of development ($\bar{L}_{\mathrm{max}}$), this was not the case for the mean group payoff ($\bar{Z}_{\mathrm{max}}$) and the number of innovated traits ($\bar{T}_{\mathrm{max}}$), which peak at intermediate values of $k$ and monotonously decrease with $k$, respectively. These non-trivial behaviors indicate that cultural accumulation is subject to an exploration-exploitation trade-off between innovation (exploration) and social learning (exploitation), a phenomenon that was previously also observed in both experimental and theoretical works on social dynamics including cultural evolution [14, 33], wisdom of the crowds [45], as well as collective problem-solving [46, 36, 47, 48].
Dense random networks (i.e., networks with high values of $k$) generally allow for the efficient dissemination of information about cultural traits (exploitation) and therefore lead to high cultural complexity. However, this comes at the expense of cultural diversity (number of traits), as the rate of innovation is impaired due to an over-reliance on social learning. Furthermore, previous empirical and theoretical findings in cultural evolution and social learning studies suggest that collective performance depends on both the complexity of the task and the underlying network structure. In particular, it was found that performance in dense networks is often decreased for complex tasks [36, 47, 48], a phenomenon best explained by sufficient levels of transient diversity in less densely connected systems [49]. In Fig. 3, the shapes of $\bar{L}_{\mathrm{max}}$ and $\bar{Z}_{\mathrm{max}}$, as functions of $k$, therefore indicate that the different sub-processes of cultural accumulation vary in complexity: While reaching a high level of complexity is a rather simple problem (high value of $\bar{L}_{\mathrm{max}}$ in dense networks), it is more complex to collectively coordinate on high average payoffs (low values of $\bar{Z}_{\mathrm{max}}$ in dense networks). In other words, populations develop traits and reach high complexity rather easily, however, such traits do not necessarily have a high utility (i.e. high payoff). In our model, collective payoffs are optimal at a medium level of connectivity, where agents optimally solve the explore-exploit trade-off, i.e., have access to the right amount of information through social learning while maintaining sufficient incentives to innovate.
@@
In the second part, we investigated the combined impact of network-based social learning and algorithmic mediation on cultural accumulation. For a fixed value of $k$, we found that increasing $r$—the ratio between algorithmic mediation over network-based social learning—strongly affects the overall dynamics: while the cultural complexity monotonously increases with $r$, the mean payoff gives rise to maximum values for intermediate levels of algorithmic mediation, which suggests a non-trivial interaction between $k$ and $r$. The less we are connected, the more we can benefit from algorithmic mediation. For a given link density, however, the benefits of increased information sharing via algorithmic mediation become outweighed by their detrimental consequences: too high levels of $r$ create too much effective connectivity, which decreases performance as $r\rightarrow 1$. This result is in line with the strongly decreasing number of traits as $r$ increases, indicating low levels of cultural diversity. Interestingly, this behavior remains remarkably stable when the structure of the underlying network is altered, provided the link density is kept constant, as demonstrated in the Supplemental Material (SM). For two additional network structures—small-world networks [50] and scale-free networks [51]—the mean group level, the number of traits, and the payoffs exhibit the same qualitative behavior as observed in random ER networks.
Furthermore, we examined alternative social learning strategies (in addition to the payoff-biased learning addressed previously) and their influence on cultural accumulation in the presence of algorithmic mediation. Two strategies were considered: one in which cultural traits are copied from the neighbour with the highest complexity level, and alternatively, one in which a random neighbour is chosen. The latter strategy is analogous to the unbiased transmission described in [31]. It can be seen that for $r=0$ (solely social learning), the random strategy performs considerably less well than the other two, but the situation becomes much more intricate when algorithmic mediation is introduced (see Figure 2 in SM). The mean number of levels as a function of $r$ is considerably lower for the random selection strategy, yet it yields larger payoffs for high values of $r$. This seemingly contradictory effect is the result of two mechanisms: firstly, a randomly selected neighbour is rarely at a sufficiently high level to allow for copying, which leads to a high number of innovations (as evidenced by the values of the mean number of traits in the middle panel of Figure 2 in SM); secondly, when the proportion of algorithmic recommendations is sufficiently high, the innovated items are efficiently distributed throughout the system. Consequently, a seemingly not effective random strategy at the individual level becomes optimal for the population.
These results make a clear case for studying the dynamics of algorithmic mediation not in isolation but in relation to social learning strategies motivated by cognitive science (e.g., payoff-biased learning). In our model, both algorithmic mediation and social learning acting alone consistently lead to worse payoffs and fewer cultural traits than a combination of both processes. This behavior may point toward a general trade-off in socio-technical systems between access to information (and thus the possibility to exchange it) and the ability of individuals to innovate: while algorithms can be beneficial for cultural accumulation, they may also lead to detrimental effects if used excessively [52, 53].
The real-world counterparts of these dynamics provide valuable insights into modern socio-technical systems. For example, recommender systems and large language models (LLMs) shape cultural accumulation by redistributing information globally, facilitating knowledge acquisition and skill development. However, they also risk suppressing diversity or incentivizing conformity when over-relied upon. Applications in collaborative platforms [54, 55], professional networks [56], or online communities [30, 57] could serve as real-world analogs of the processes modeled here. Empirical studies might examine how varying levels of algorithmic support in such environments affect innovation rates or cultural diversity, shedding light on the interplay between cultural exploration and exploitation in digital ecosystems [58, 16, 59, 60].
Our model has several limitations that should be addressed in future work. First, it focuses exclusively on algorithmic mediation, i.e. the redistribution of existing cultural traits and does not account for algorithmic innovation, where intelligent systems actively generate new solutions or artifacts. Previous research has found strong support for the ability of machines to also take part in the innovation process [61, 62], which future models of cultural evolution should take into account. Second, we assume that algorithmic mediation remains static, whereas the behavior of the real-world algorithms co-evolves with human behavior driven by user feedback and iterative improvement of machines [60]. Third, our model does not include biases in algorithmic mediation, such as leading to unintended distortions in content distribution, all of which are prevalent in real systems [63]. Furthermore, extending our framework to account for more complex cultural spaces will be an essential next step for capturing the richness of real-world cultural dynamics, including more complex innovation processes such as recombination [39]. Ultimately, the predictions of our model on the interaction between algorithmic mediation and social learning must stand up to empirical testing through controlled laboratory experiments and observational studies in diverse online and offline environments.
Authors’ Contributions
A.C.: conceptualization, investigation, software, visualization, writing—original draft, writing—review and editing; F.B.: conceptualization, investigation, visualization, writing—original draft, writing—review and editing; I.R.: conceptualization, investigation, writing—original draft, writing—review and editing.
All authors gave final approval for publication and agreed to be held accountable for the work performed therein.
Acknowledgments
The authors would like to thank the reviewers for their constructive comments.
Data Accessibility
Simulation code to reproduce the results in this paper is publicly available in GitHub repository [64] and permanently archived on Zenodo [65]. Electronic supplementary material is available online [66].
Funding Statement
No additional funding was received for this study.